Don runs a charity fruit sale, selling boxes of oranges for $11 and boxes of grapefruit for $10. If he sold a total of 762 boxes and took in $8125 in all, then how many boxes of oranges did he sell?
Answer by Nate(3500)

Don runs a charity fruit sale, selling boxes of oranges for $11 and boxes of grapefruit for $10. If he sold a total of 762 boxes and took in $8125 in all, then how many boxes of oranges did he sell?
# of Boxes of Oranges = x
# of Boxes of Grapefruit = y
x + y = 762 or x = 762 - y
Plug:
11x + 10y = 8125
11(762 - y) + 10y = 8125
8382 - 11y + 10y = 8125
257 = y
x = 762 - y = 762 - 257 = 505
257 boxes of grapefruit
505 boxes of oranges